# Graph Neural Network-Based Reinforcement Learning for Controlling Biological Networks: The GATTACA Framework

### v1

@@ 2.6.1 Graph Convolutional Operation
In our implementation, a graph convolutional layer takes the values of each input node from the previous layer and produces the corresponding output vectors for each node. The convolutional layer implements the function:
@@
where $v$ is the current node, $N(v)$ is the set of neighbours of $v$, and $h$ is the graph convolutional kernel with learnable parameters $\Theta$. In our case $h$ consists of a neural network with a 64-neuron deep layer and ReLu activations. Here $h(x||y)$ means conditioning-by-intervention, that is the distribution of $x$ with $y$ set to some fixed value. It’s role is to reduce spurious correlations present in the graph; for more details, we refer to [35]. Our network consists of three graph convolutional layers. The output of graph convolutions is then passed through a three-hidden-layer MLP with 1024, 512, and 256 neurons per layer. Each of the attention branches, as well as the state value layer, is an MLP with (256, 512, 512) neurons. We expect that our network will learn interactions between genes, and create an embedding of a discreet graph into continuos space, which is more suitable for machine learning algorithms.
3 Control problem formulation
We proceed to formulate a novel version of the target control problem tailored for cellular reprogramming, understood as an artificial change of the identity or functional state of a cell by manipulating its gene expression profile.
Often biologists are interested not in a single state but a set of states in which a specified subset of genes attains certain configuration of expression levels which afterwards remains constant. For example, they may be interested in achieving any state in which apoptosis occurs. Therefore, we assume that the target is defined as a combination of desired values for a subset of genes. With this, we can proceed with the definition of an attractor-target control strategy.
source attractor state target1target2target3controlcontrolcontroltarget basinintermediate basinintermediate basin target configuration
Figure 1: Schematic illustration of the concept behind Attractor-Target Control. The target state can be any of the attractor states that are aligned with the target configuration. Transitions that follow the original network dynamics are shown as dashed arrows. Red lightnings indicate interventions. The depicted control strategy is of length 3.
Definition 3.1 (Attractor-Target Control Strategy).
Given a BN, a source (pseudo-)attractor state, and a specified target configuration for a subset of genes, an attractor-target control strategy is a sequence of interventions that guides the BN dynamics from the source (pseudo-)attractor state to some (pseudo-)attractor state that aligns with the desired gene configuration. Interventions are understood as simultaneous flips (perturbations) of values for a subset of genes in a particular network state and their application is limited to (pseudo-)attractor states.
Furthermore, length of a control strategy is defined as the number of interventions in the control sequence. We refer to a target control strategy of the shortest length as the minimal attractor-target control strategy.
The concept of the attractor-target control strategy is schematically illustrated in Figure 1. Notice that, according to the above definition, we aim to guide the network dynamics toward any state within a subset of (pseudo-)attractor states that match the target configuration, rather than a specific one. This definition is more general than the definition of the ‘Attractor-Based Control Strategy’ introduced in [17], where the system is expected to reach a specific attractor. The current definition of the ‘Attractor-Target Control Strategy’ aligns better with contemporary biological practice.
Definition 3.2 (Attractor-Target Control).
4 Graph-based Attractor-Target Control Algorithm
We propose herein a Graph-based Attractor-Target Control Algorithm (GATTACA) – a GNN-based DRL framework for the attractor-target control problem. We focus on cellular phenotypic functional states, which are observable in experimental practice, and thus allow the DRL agent, which we refer to as the Attractor-Target Graph-based Controller (ATGC), to take actions only in (pseudo-)attractor states in agreement with our control problem formulation in Section 3. Since the identification of all attractor states is an NP-hard problem in itself, we further develop the pseudo-attractor identification procedure (PASIP), which we previously established in [17]. The source of GATTACA and the trained models are available at [37].
4.1 Improved PASIP
We observed that the original PASIP method in [17] can be impeded by the presence of large attractors in a BN model. For example, the T-LGL network of [38] has an attractor with more than 1000 states. Of course, if the existence of a large attractor is foreseen and a rough estimate of its size is available, one could adjust the PASIP hyperparameters appropriately. However, this would imply the necessity of longer simulations of the BN model, which are computationally expensive. If the existence of a large attractor is not anticipated and the hyper-parameters are not adjusted, the original PASIP is prone to getting trapped in such an attractor while waiting for one of the attractor states to demonstrate a pronounced frequency of re-visits. To handle such cases effectively, we modify PASIP by adding additional checkpoints that enable the identification of a representative state within a large attractor. The improved PASIP procedure (iPASIP) consists of steps executed in two phases: (I) environment pre-processing and (II) DRL agent training. It extends the original PASIP from [17] by incorporating Steps I-2 and II-3 as the additional checkpoints.
  1. I-1
    During environment pre-processing, $k$ PBN simulations are started from randomly selected initial states. Each PBN simulation runs for initial $n_{0}=200$ time steps, which are discarded as the so-called burn-in period. Then, the simulation continues for $n_{1}=1000$ time steps during which the visits to individual states are counted. All states in which at least $5\%$ of the simulation time $n_{1}$ is spent are added to the list of pseudo-attractor states.
  2. I-2
    If at any point more than 1,000,000 steps are taken without encountering a dominating state (one revisited at least 5% of the time), it is assumed that the dynamics have become trapped in a large attractor. The most frequently encountered state is then added to the list of pseudo-attractor states as a representative of this plausible attractor.
  3. II-1
    The simulation of the PBN environment may enter a fix-point attractor not detected in Step I. If the simulation gets stuck in a particular state for $n_{2}=1000$ steps, the state is added to the list of pseudo-attractor states.
  4. II-2
    During training, the simulation of the PBN environment may enter a multi-state attractor that was not detected in Step I. To account for this, a history of the most recently visited states is maintained. When the history buffer reaches a size of $n_{3}=10000$ items, revisits to each state are counted and any state revisited more than 15% of the time is added to the list of pseudo-attractor states. If no such state exists, the history buffer is cleared and the procedure continues for another $n_{3}$ time steps. Otherwise, the history information is discarded.
  5. II-3
    As in I-2, if at any point more than 1,000,000 steps are taken without encountering a dominating state (one revisited at least 5% of the time), it is assumed that the dynamics have become trapped in a large attractor and a representative of this plausible attractor is added to the list of pseudo-attractor states.
We notice that while the new checkpoints may clearly introduce false-positive pseudo-attractor states, this occurs rarely. Importantly, the representative states of plausible large attractors stand out due to being frequently revisited. Even if these states are not exact pseudo-attractor states, to a large extent they resemble them. Therefore, given their rarity, adding them to the list of pseudo-attractor states is acceptable. This is confirmed by the results of our case studies, presented and discussed in Section 6.
Our control learning algorithm can handle cases where the number of pseudo-attractors is greater than 25,000. Moreover, we conducted experiments in which the number of pseudo-attractor states was overestimated by a factor of 1000 and ATGC was still able to identify effective control strategies (data not shown). Those two facts prove that the ATGC is resilient to both high attractor counts and spurious pseudo-attractor states.
Since iPASIP method can introduce spurious pseudo-attractor states, we will henceforth use the term ‘pseudo-attractor state’ in a broader sense: it will refer to all states returned by iPASIP rather than only those defined in Definition 2.1.
4.2 Attractor-Target Graph-based Controller (ATGC)
@@
In our training framework, we specify a source pseudo-attractor state and a target configuration that specifies the desired values for a subset of genes. Then, we let the DRL agent perform actions, i.e. interventions consisting of simultaneous perturbation of up to five genes once a pseudo-attractor state different from the target is reached. After an action, we let the environment evolve in accordance with its original dynamics into the next pseudo-attractor state. This models the biological process of transforming a cell from one phenotypic functional state into another by conducting a series of in vivo reprogramming interventions.
Figure 2: Schematic representation of the ATGC DRL agent of GATTACA.
We define a new with respect to the one in [17] DRL reward function as
@@
where $s$ and $s^{\prime}$ are the current and next pseudo-attractor states, respectively, $\mathbbm{1}_{TA}$ is an indicator function of the target configuration, and $|a|$ is the number of genes perturbed by applying action $a$.
This reward scheme is balanced to punish the model for perturbing many genes in a single action while providing enough reward to allow to still reach the target configuration by the means of a sequence of actions. Our previous experiments showed that it is important that the reward function maintains a consistent sign during training (in this particular case it is designed to remain non-negative, which is ensured by the term of 21). For the loss function, We use the Double Deep Q Network (DDQN) loss function as in [18] and [17], that is the Mean Squared Error (MSE) between the predicted Q-values and the target Q-values, calculated using equation:
$\mathcal{L}(\theta)=\mathbb{E}\left[\left(q_{\pi}(s,a)-target(s,a)\right)^{2}\right],$
where $target$ is the delayed copy of the $q_{\pi}$ network.
Our DRL agent is trained in episodes. An episode is a sequence of states, actions, and rewards that starts in the specified source state and ends in any of the pseudo-attractor states that comply with the target configuration. It represents a complete trajectory of the agent’s interaction with the environment.
For each episode, we randomly choose a source (pseudo-)attractor state. We make sure that this state fulfils the imposed initial conditions and does not satisfy the target configuration. Then we let our agent apply its strategy and terminate each episode after 100 unsuccessful actions. If 100 actions are not enough to reach the target configuration, the respective episode is considered unsuccessful and terminated.
5 Experiments
We evaluate how GATTACA scales with the size of the BN model. We focus on large, verified models taken from the literature for well-known real-world biological networks.
5.1 Case Studies: Models and Their Target Configurations
To evaluate the performance of our approach, we select a variety of models from the literature, ranging in size from 35 nodes (the Bladder model) to 188 nodes (the CD4+ model). To ensure that the target configurations used in our evaluation are both realistic and of significance to biologists, we focus on the configurations considered in the original publications where these models were introduced. The models and their corresponding target configurations are listed in Table 1. Below, we briefly present each of them.
@@ 5.1 Case Studies: Models and Their Target Configurations
Table 1: BN models of real-world biological networks with their target configurations used for the performance evaluation of GATTACA. The BD Id column provides model identification numbers in the BioDivine repository [25].
@@
Table 2: Environmental conditions for the BN models of real-world biological networks used in the evaluation of GATTACA. Unspecified node values are indicated with *.
@@
5.2 Performance Evaluation Methodology
We evaluated the performance of GATTACA in solving the attractor-target control problem formulated in Section 3 on asynchronous BN models of real biological systems in a wide range of sizes using case studies described in Section 5.1.
@@ 5.2 Performance Evaluation Methodology
Since we want ATGC of GATTACA to learn the relationship between the target genes and other genes, we let it perturb all nodes (including input nodes) except the ones specified in the target configuration, which we refer to as target genes. Beyond this, we do not impose any other restrictions on genes that can be perturbed: we allow full flexibility in selecting genes for interventions to ensure a fair comparison with the CABEAN benchmark and to demonstrate the potential of GATTACA to identify optimal or suboptimal strategies in large networks. However, we emphasize that our framework can be easily modified to incorporate previous biological knowledge by restricting perturbations to a predefined subset of genes. In our experiments, we specify the target genes and their configurations based on the literature from which the models were sourced. The details are presented in Section 5.1.
@@
Furthermore, the more genes that need to be perturbed in a single intervention, the more challenging and expensive the experimental procedure becomes. As a result, the in vitro implementation of the predicted control strategy may be unfeasible. Hence, to make our approach as cost-effective as possible, we decide to limit the number of nodes that can be perturbed simultaneously in a single intervention. This number is a parameter of our method and we set it to five in our experiments. This setting suffices to compute successful control strategies for all of our case studies. Nevertheless, the parameter value can be adjusted to particular needs as required.
@@
ATGC is trained to minimise the number of control actions required to drive a BN model from a randomly chosen source attractor to any attractor in which the given target configuration is realised. Specifically, for a specified environmental condition, a training episode begins by randomly selecting one of the pseudo-attractor states identified so far by iPASIP that does not align with the given target configuration. Throughout each episode, we track the number of control actions taken, referred to as the episode control length, as ATGC guides the network dynamics toward any pseudo-attractor state aligned with the given target configuration. ATGC can take up to 100 control actions per episode. If it fails to control the BN model within this limit, the episode is considered unsuccessful, indicating that no valid control strategy was found.
@@
To evaluate the performance of GATTACA on a particular BN, we recover control strategies by simulating the environment from a randomly selected source pseudo-attractor state that is not aligned with the target configuration. At each intermediate pseudo-attractor state, a control action predicted by the trained DRL agent is applied. For each recovered control strategy, we record its length and the information whether the target configuration is reached. Due to the nondeterministic dynamics of the environment introduced by the asynchronous update mode, results may vary between runs. To address this issue, we recover a control strategy 10 times for each of the randomly selected source pseudo-attractor states. For a given BN model, we report the average length of successful control strategies.
@@
We compare the results of GATTACA with those obtained using the state-of-the-art exact attractor-based sequential instantaneous source-target control (ASI) algorithm introduced in [16] and implemented in the CABEAN software tool [30]. The exact algorithm follows the divide-and-conquer approach based on SCC-decomposition of BN structure graph and it leverages symbolic techniques based on Binary Decision Diagrams (BDDs). Amongst the control algorithms provided by CABEAN, we selected ASI as the best match for the control problem in this study.
@@
For a fair comparison, we set in CABEAN the initial values of the input nodes of a BN model in accordance with a particular environmental condition. CABEAN requires at least some of the input nodes to have fixed values. Therefore, in the case of partially specified environmental conditions, i.e. for models Bladder, MAPK, and CD4+, it is possible to consider all of the corresponding fully specified environmental conditions at once as is the case for GATTACA.
@@
However, in the case of all environmental conditions considered, e.g. the T-LGL model, fully specified environmental conditions are enumerated and considered one by one in CABEAN. For each such condition, all attractors are computed with CABEAN and divided into source attractors, i.e. those not aligned with the target configuration, and target attractors, i.e. those aligned with the target configuration. Then, for each source attractor, the shortest path to any of the target attractors is computed with CABEAN. If there is no available target attractor for a given environmental condition, we check all other environmental conditions for a target attractor. We first switch the environmental condition, and then (if necessary) use CABEAN to control the network within the new environmental condition. The environmental condition switch adds 1 to the control path length in CABEAN. We take the minimum length over all other environmental conditions containing a target attractor. We provide an example of this scenario in Appendix B. This behaviour is mimicked by GATTACA by letting it perturb input nodes. Finally, the averaged length of the shortest paths over all environmental conditions and source attractors is reported. Formally, the average control path length determined by CABEAN for a BN model $M$, denoted $ACPL_{CABEAN}(M)$, id defined as follows:
@@
$\text{ACPL}_{\text{CABEAN}}(M)\stackrel{{\scriptstyle\text{def}}}{{=}}\mathbb{E}_{e\in E_{M}}\left[\mathbb{E}_{s\in\mathcal{S}^{M}_{e}}\left[\begin{cases}\min\limits_{t\in\mathcal{T}^{M}_{e}}\ \text{cpl}(s,t)&\text{if }\mathcal{T}^{M}_{e}\neq\emptyset\\ \min\limits_{e^{\prime}\in E_{M}}\left(1+\min\limits_{t\in\mathcal{T}^{M}_{e^{\prime}}}\ \text{cpl}(s,t)\right)&\text{otherwise}\end{cases}\right]\right],$ (3)
@@
where $E_{M}$ is the set of all environmental conditions considered for model $M$, $\mathcal{S}^{M}_{e}$ is the set of source attractors and $\mathcal{T}^{M}_{e}$ is the set of target attractors for a given environmental condition $e$ for model $M$, and cpl is the control path length computed by CABEAN for a specific pair of source and target attractors. Notice that in the case of a partially specified environmental condition where all the corresponding fully specified conditions can be handled by CABEAN at once, the set $E_{M}$ in Equation 3 is a singleton. In the case where all environmental conditions are considered, the cardinality of $E_{M}$ is equal to the number of distinct possible settings of the values for all input nodes.
6 Results

### v2

@@ V-B Attractor-Target Graph-based Controller (ATGC)
The GCN consists of three graph convolutional layers. In our implementation, a graph convolutional layer takes the values of each input node from the previous layer and produces the corresponding output vectors for each node. The convolutional layer implements the function:
@@
where $v$ is the current node, $N(v)$ is the set of neighbours of $v$, and $h$ is the graph convolutional kernel with learnable parameters $\Theta$. In our case, $h$ consists of a neural network with a 64-neuron deep layer and ReLu activations. Here $h(x||y)$ means conditioning-by-intervention, that is the distribution of $x$ with $y$ set to some fixed value. Its role is to reduce spurious correlations present in the graph; for more details, we refer to [28].
The reason for introducing the GCN into the architecture of the ATGC is twofold. First, to allow the DRL agent to effectively learn the interactions between the BN nodes that represent the system components. In particular, the available knowledge of the structure of the BN is explicitly incorporated into the ATGC by defining the GCN architecture in accordance with the structure graph of the BN. Then, during training, the BN is simulated in accordance with its Boolean functions. Since the resulting state of the BN is observed and used to compute the reward function discussed in the following, the information on the BN dynamics is implicitly encoded via training into the graph convolutional kernels $h_{\Theta}$ and the remaining parameters of the ATGC architecture.
Second, the GCN creates an embedding of a discrete graph into a continuous space, which is more suitable for machine learning algorithms and eases the training of the DRL agent.
The output of the graph convolutions is passed through a three-hidden-layers MLP with 1024, 512, and 256 neurons per layer.
Finally, to allow simultaneous perturbation of a combination of genes within a DRL action as in [7], the MLP is followed by a BDQ network architecture [29] [^1] An architecture consisting of the MLP followed by a DDQN was originally proposed in [30]. In [7], we replaced the DDQN with the BDQ architecture to allow multiple perturbations of genes within a single DRL action. For further details, refer to [7].. Each of the five action branches and the state value layer in Figure 2 is a three-layer MLP with 256, 512, and 512 neurons. The idea of including the state value layer is to create the internal representation of the current state, so that it can later be included in the calculation of the Q function of the state-action pair. We want all of the branches to share this common representation, as the value of each action changes depending on the state. We consider five action branches as this is the maximum number of nodes in our study that can be perturbed simultaneously within a single DRL action.
@@
Training in the GATTACA framework is performed as follows. We specify a source pseudo-attractor state and a target configuration that determines the desired values for a subset of genes. We then let ATGC perform actions, i.e. interventions consisting of simultaneous perturbation of up to five genes in a PA state that is not aligned with the target configuration. After an action, we let the environment evolve in accordance with its original dynamics into the next PA state. This models the biological process of transforming a cell from one phenotypic functional state into another by conducting a series of in vivo reprogramming interventions.
Figure 2: Schematic representation of the ATGC DRL agent of the GATTACA framework.
We define the DRL reward function as follows:
@@
where $s$ and $s^{\prime}$ are the current and next PA states, respectively, $\mathbbm{1}_{TA}$ is an indicator function of the target configuration, and $|a|$ is the number of genes perturbed by applying action $a$.
This reward scheme is balanced to punish the model for perturbing many genes in a single action while providing enough reward to allow to still reach the target configuration by the means of a sequence of actions. Our previous experiments showed that it is important that the reward function maintains a consistent sign during training (in this particular case, it is designed to remain non-negative, which is ensured by the term of 21). For the loss function, we use the $\mathcal{L}_{t}^{\textrm{BDQ}}$ function defined in Section III-E2.
Our DRL agent is trained in episodes. An episode is a sequence of states, actions, and rewards that starts in the specified source state and ends in any of the PA states that comply with the target configuration. It represents a complete trajectory (path) of the agent’s interaction with the environment.
For each episode, we randomly choose a source PA state. We make sure that this state does not satisfy the target configuration. Then we let our agent apply its strategy. If applying 100 actions within an episode is not sufficient to reach the target configuration, the respective episode is considered unsuccessful and terminated.
VI Experiments
We evaluate how our GATTACA framework scales with the size of the BN model. We focus on large, verified models taken from the literature for well-known real-world biological networks.
VI-A Case Studies: Models and Their Target Configurations
To evaluate the performance of our approach, we select a variety of models from the literature, ranging in size from 35 nodes (the Bladder model) to 188 nodes (the CD4+ model). To ensure that the target configurations used in our evaluation are both realistic and of significance to biologists, we focus on the configurations considered in the original publications where these models were introduced. The models and their corresponding target configurations are listed in Table I. Below, we briefly present each of the models.
@@ VI-A Case Studies: Models and Their Target Configurations
TABLE I: BN models of real-world biological networks with their target configurations used for the performance evaluation of the GATTACA framework. The BD Id column provides model identification numbers in the BioDivine repository [31].
@@
TABLE II: Environmental conditions for the BN models of real-world biological networks used in the evaluation of GATTACA. Unspecified node values are indicated with *.
@@
VI-B Performance Evaluation Methodology
We evaluated the performance of the GATTACA framework in solving the attractor-target control problem formulated in Section IV on asynchronous BN models of real biological systems in a wide range of sizes using case studies described in Section VI-A.
@@ VI-B Performance Evaluation Methodology
Since we want ATGC of our GATTACA framework to learn the relationship between the target genes and other genes, we let it perturb all nodes (including input nodes) except the ones specified in the target configuration, which we refer to as target genes. Beyond this, we do not impose any other restrictions on genes that can be perturbed: we allow full flexibility in selecting genes for interventions to ensure a fair comparison with the CABEAN benchmark and to demonstrate the potential of the GATTACA framework to identify optimal or suboptimal strategies in large networks. However, we emphasize that our framework can be straightforwardly modified to incorporate prior biological knowledge by restricting perturbations to a predefined subset of genes.
@@
Furthermore, the more genes that need to be perturbed in a single intervention, the more challenging and expensive the wet-lab experimental procedure becomes. As a result, the in vitro implementation of the control strategy predicted by the GATTACA framework may be unfeasible. Hence, to make our approach as cost-effective as possible, we decide to limit the number of nodes that can be perturbed simultaneously in a single intervention. This number is a parameter of our framework and we set it to five in our experiments. This setting suffices to compute successful control strategies for all of our case studies. Nevertheless, the parameter value can be adjusted to particular needs as required.
@@
ATGC is trained to minimise the number of control actions required to drive a BN model from a randomly chosen source attractor to any attractor in which the given target configuration is realised. Specifically, for a specified environmental condition, a training episode begins by randomly selecting one of the PA states identified so far by iPASIP that does not align with the given target configuration. Throughout each episode, we track the number of control actions taken, referred to as the episode control length, as ATGC guides the network dynamics toward any PA state aligned with the given target configuration. ATGC can take up to 100 control actions per episode. If it fails to control the BN model within this limit, the episode is considered unsuccessful, indicating that no valid control strategy was found.
@@
To evaluate the performance of the GATTACA framework on a particular BN, we recover control strategies by simulating the environment from a randomly selected source PA state that is not aligned with the target configuration. At each intermediate PA state, a control action predicted by trained ATGC is applied. For each recovered control strategy, we record its length and the information whether the target configuration is reached. Due to the nondeterministic dynamics of the environment introduced by the asynchronous update mode, results may vary between runs. To address this issue, we recover a control strategy 10 times for each of the randomly selected source PA states. For a given BN model, we report the average length of successful control strategies.
@@
We compare the results of the GATTACA framework with those obtained using the state-of-the-art exact attractor-based sequential instantaneous source-target control (ASI) algorithm introduced in [39] and implemented in the CABEAN software tool [5]. The exact algorithm follows the divide-and-conquer approach based on SCC-decomposition of BN structure graph and it leverages symbolic techniques based on Binary Decision Diagrams (BDDs) to encode the BN dynamics. Amongst the control algorithms provided by CABEAN, we selected ASI as the best match for the control problem considered in this study.
@@
For a fair comparison, we set in CABEAN the initial values of the input nodes of a BN model in accordance with a particular environmental condition. CABEAN requires at least some of the input nodes to have fixed values. In the case of partially specified environmental conditions of Bladder, MAPK, and CD4+ models it is possible to consider in CABEAN all of the corresponding fully specified environmental conditions at once as is the case of our framework.
@@
However, in the case of all environmental conditions considered, e.g. the T-LGL model, fully specified environmental conditions are enumerated and considered one by one in CABEAN. For each such condition, all attractors are computed with CABEAN and divided into source attractors, i.e. those not aligned with the target configuration, and target attractors, i.e. those aligned with the target configuration. Then, for each source attractor, the shortest path to any of the target attractors is computed with CABEAN. If there is no available target attractor for a given environmental condition, we check all other environmental conditions for a target attractor. We first switch the environmental condition, and then (if necessary) use CABEAN to control the network within the new environmental condition (the GATTACA framework can switch environmental conditions automatically by letting it perturb input nodes). We provide an example of this scenario in Appendix -B. To account for the “manual” environmental condition switch, one is added to the control path length returned by CABEAN. We take the minimum length over all other environmental conditions containing a target attractor. Finally, the averaged length of the shortest paths over all environmental conditions and source attractors is reported.
Formally, the average control path length determined by CABEAN for a BN model $M$, denoted $ACPL_{CABEAN}(M)$, is defined as follows:
{strip}
@@
$\text{ACPL}_{\mathrm{CABEAN}}(M)\;=\;\mathbb{E}_{e\in E_{M}}\Biggl{[}\,\mathbb{E}_{s\in\mathcal{S}^{M}_{e}}\Biggl{[}\begin{cases}\displaystyle\min_{t\in\mathcal{T}^{M}_{e}}\!\text{cpl}(s,t)&\text{if }\mathcal{T}^{M}_{e}\neq\emptyset,\\[4.30554pt] \displaystyle\min_{e^{\prime}\in E_{M}/\{e\}}\Bigl{(}1+\min_{t\in\mathcal{T}^{M}_{e^{\prime}}}\!\text{cpl}(s,t)\Bigr{)}&\text{otherwise.}\end{cases}\Biggr{]}\Biggr{]},$ (2)
@@
where $E_{M}$ is the set of all separate environmental conditions considered for model $M$, $\mathcal{S}^{M}_{e}$ is the set of source attractors, $\mathcal{T}^{M}_{e}$ is the set of target attractors for a given environmental condition $e$ for model $M$, and cpl is the control path length computed by CABEAN for a specific pair of source and target attractors. Notice that in the case of a fully or partially specified environmental condition where all the corresponding fully specified conditions can be handled by CABEAN at once, the set $E_{M}$ in Equation (2) is a singleton. In the case where all environmental conditions are considered, the cardinality of $E_{M}$ is equal to the number of distinct possible settings of the values for all input nodes.
VII Results